In a purse is $3.05 in nickels and dimes, 19 more nickels than dimes. How many coins are there of each kind?
Answer by checkley77(12844)

N=D+19
.05(D+19)+.10D=3.05
.05D+.95+.10D=3.05
.05D+.10N=3.05-.95
.15D=3.05-.95
.15D=2.10
D=2.10/.15
D=14 NICKELS.
N=14+19
N=33 DIMES.
PROOF:
.05*33+.10*14=3.05
1.65+1.40=3.05
3.05=3.05